### v1

@@ 3.1.2 Weak formulation of the Steklov problem
The following theorem follows from the characterisation [AretE15, Proposition 5.1] of the operator domain $\operatorname{Dom}\left(A^{\mathrm{D}}\right)$ and [AuHa14b, Section 9].
@@
Additionally, the corresponding eigenfunctions $u_{k}$ can be chosen to form a complete orthonormal basis of the subspace $\left\{u\in E^{1}\left(\Omega^{\mathrm{ext}}\right):\Delta u=0\right\}$ and the restrictions of these eigenfunctions to $\partial\Omega$ form a complete orthonormal basis of $L^{2}(\partial\Omega)$.
@@
Arendt and ter Elst [AretE15] proposed two ways to approximate the Dirichlet-to-Neumann operator on exterior domains by truncated domains. For any $R>R_{0}(\Omega)$, recalling that $\Omega^{\mathrm{ext}}_{R}=\Omega^{\mathrm{ext}}\cap B_{R}$, we consider the Dirichlet approximation,
@@
$\begin{cases}\Delta u_{R}=0\qquad&\text{ in }\Omega^{\mathrm{ext}}_{R},\\ \partial_{\nu}u_{R}=\sigma^{\mathrm{N}}u_{R}\qquad&\text{ on }\partial\Omega,\\ \partial_{\nu}u_{R}=0\qquad&\text{ on }\partial B_{R}.\end{cases}$ ($\mathrm{SN}_{R}$)
@@
For dimensions $n\geq 3$, it was already proved in [AretE15] that the truncated domain approaches are equivalent to the approaches from §3.1.1.
Theorem 3.9 ([AretE15, Theorems 5.5 and 5.6]).
@@
where $\mu=\mu\left(\Omega^{\mathrm{ext}},\Lambda\right)$ is the spectral parameter. Since $\Lambda^{2}>0$, the corresponding bilinear form is $H^{1}\left(\Omega^{\mathrm{ext}}\right)$-coercive. In [Au12, Sections 4 and 8], it is demonstrated that coercivity can be used to show that the spectrum of a Steklov problem on bounded domains is purely discrete. Similarly, by applying the same modifications to [Au12] as in, e.g., [AuHa13, Theorem 5.1], we deduce that the spectrum of ($\mathrm{H}_{\Lambda}$) is purely discrete, with the eigenvalues
$\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\min_{\begin{subarray}{c}M\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)\\ \dim(M)=k\end{subarray}}\max_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S},$ (3.8)
and with the eigenfunctions belonging to $H^{1}\left(\Omega^{\mathrm{ext}}\right)$. In addition, the eigenfunctions are all exponentially decaying as $|x|\to\infty$ due to
@@
We now consider the cases of the balls of dimension $n\geq 3$ and disks in the planar case $n=2$ separately.
@@
For $n=2$, we have the same form of the eigenfunctions as in (3.9). For $\ell\geq 1$, we act as before. For $\ell=0$, we use additionally $K_{0}(x)=-\log\frac{x}{2}+O(1)$ as $x\searrow 0$. This yields
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. We assume that the origin $0\in\Omega$, which can be always achieved by a change of variables, and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$ by interpreting $x=\left(x_{1},x_{2}\right)\in\mathbb{R}^{2}$ as $z=x_{1}+\mathrm{i}x_{2}\in\mathbb{C}$.
@@
We will show later that (3.11) allows us to adapt classical arguments by Weinstock to derive isoperimetric-type inequalities for the first non-trivial Steklov eigenvalue. First, however, we establish the equivalence between the formulation ($\mathrm{CT}$) and the approaches described in §3.2 and §3.3.
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$ is a bounded open set with a Lipschitz boundary $\partial\Omega$. Let $\Phi$ be the fundamental solution for the Laplace equation,
@@
Finally, let us recall that the single layer operator is always invertible in dimension $n=3$, whereas for $n=2$ we have to assume a suitable scaling of the domain $\Omega$ [St07, Section 6.6]. If $n=2$ and the logarithmic capacity of $\Omega$ is equal to one, the single layer potential has a nontrivial kernel and one cannot use (4.1) to define the eigenvalues of the Dirichlet-to-Neumann operator.
@@
While the boundary integral equation (4.1) remains valid for characterising those eigenfunctions whose harmonic extensions vanish at infinity (up to switching the sign), see §4.3.1, the situation differs in dimension two: as noted in §2.1, exterior Steklov eigenfunctions in two dimensions are only required to remain bounded and do not need to decay. The Dirichlet-to-Neumann operator in exterior domains, including dimension two, together with its Fredholm properties, has been analysed in [KacNaPa25, Section 3]. Similarly, we obtain a representation of the Dirichlet-to-Neumann operator in terms of boundary layer potentials. In contrast to [KacNaPa25], our formulation also incorporates the limit at infinity via layer potentials, yielding a representation that is particularly suitable for numerical computation of eigenvalues.
@@
Since $u$ is harmonic, Green’s identities on $\Omega^{\mathrm{ext}}_{R}$ yield
@@
Note that also $\operatorname{V}_{0}u$ is well defined for $u\in H^{1}_{\text{loc}}(\Omega^{\mathrm{ext}})$, because the restriction of $u$ to $\partial\Omega$ is in $H^{\frac{1}{2}}(\partial\Omega)$ and $H^{\frac{1}{2}}(\partial\Omega)\hookrightarrow H^{-\frac{1}{2}}(\partial\Omega)$ continuously. Additionally, we define the functional $\operatorname{K}_{0}:H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$, as
@@
Accordingly, the boundary condition, $\partial_{\nu}u=\tau u$ on $\partial\Omega$ is equivalent to
$\frac{u}{2}-\operatorname{K}_{0}(u)+(\operatorname{K}u)=\tau(\operatorname{V}_{0}u)\quad\text{ on }\partial\Omega.$ ($\mathrm{LP}_{2}$)
@@
where $\left(\cdot,\cdot\right)_{\partial\Omega}$ means the duality pairing $H^{-\frac{1}{2}}(\partial\Omega)\times H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$. Additionally, let $w_{\mathrm{eq}}\in H^{-\frac{1}{2}}(\partial\Omega)$ be the unique function satisfying
@@
$w=w_{*}+\alpha w_{\mathrm{eq}},\qquad\text{with}\quad w_{*}\in H_{*}^{-\frac{1}{2}}(\partial\Omega),\quad\alpha=\left(w,1\right)_{L^{2}(\partial\Omega)}\in\mathbb{R}.$
@@
$\operatorname{W}(w_{\mathrm{eq}})=\operatorname{V}(w_{\mathrm{eq}})(0)+\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}\left(w_{\mathrm{eq}},1\right)_{\partial\Omega}=-\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}-\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}=0.$
Thus, if $\operatorname{V}_{0}(\alpha w_{\mathrm{eq}})=0$, then $\operatorname{V}(\alpha w_{\mathrm{eq}})=0$ and therefore $\alpha=0$ because $\operatorname{V}$ is invertible. ∎
@@ 4.3.1 Decay at infinity
$\frac{1}{2}u+\operatorname{K}u=\tau\operatorname{V}u$ ($\mathrm{LP}_{n}$)
@@
Since $\Phi(x,y)=\frac{1}{(n-2)|\mathbb{S}^{n-1}||x-y|^{n-2}}$, it is easy to see that $U(x)$ decays for $|x|\to\infty$.
@@ 4.3.2 Vanishing flow at infinity
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ be a bounded open set with Lipschitz boundary and $u\in H^{1}_{\text{loc}}(\Omega^{\mathrm{ext}})$ a harmonic function satisfying (4.9). First, note that (4.9) implies that $|\nabla u|\in L^{2}\left(\Omega^{\mathrm{ext}}\right)$. Hence, the average of $u$ as defined in (3.1) and denoted by $\langle u\rangle$, is finite. Define $w:=u-\langle u\rangle$. Then, by definition, $w\in W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$ (see §3.1.1) and therefore, by (3.4), we have $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. By Proposition 6.1, $w=O(|x|^{2-n})$ which shows that a higher-dimensional analogue the of far-field condition (4.3) is satisfied with $u_{\infty}:=\langle u\rangle$.
@@
$\begin{cases}\Delta g=0=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ g=-\frac{1}{\operatorname{Cap}(\Omega)}-\frac{1}{|\mathbb{S}^{n-1}|(2-n)|y|^{n-2}}\qquad&\text{ on }\partial\Omega,\\ g(y)=O\left(|y|^{1-n}\right)\qquad&\text{ as }|y|\to\infty.\end{cases}$
@@
$\begin{split}u_{\infty}=\int_{\partial\Omega}u(y)(-\partial_{\nu_{y}}G_{\mathrm{D}})(y)\,\mathrm{d}S_{y}=&\int_{\partial\Omega}-u(y)\frac{\langle y,\nu_{y}\rangle}{|\mathbb{S}^{n-1}||y|^{n}}\,\mathrm{d}S_{y}-\int_{\partial\Omega}g(y)(\partial_{\nu_{y}}u)(y)\,\mathrm{d}S_{y}\\ &-\int_{\partial B_{R}}\left(g(y)(\partial_{\nu_{y}}u)(y)-u(y)(\partial_{\nu_{y}}g)(y)\right)\,\mathrm{d}S_{y}\end{split}$
@@ 5.1.1 Finite-energy functions and Dirichlet truncation
One could use the resolvent convergence of $A_{R}^{\mathrm{D}}$, shown in [AretE15], to infer that the spectral problems ($\mathrm{ES}$) and ($\mathrm{SD}_{R}$) yield the same spectrum as $R\to\infty$. Note that the Neumann formulation ($\mathrm{SN}_{R}$) leads to a different spectrum. We instead proceed directly via the variational characterisation. This approach yields a more direct and unified framework which is also applicable to  ($\mathrm{H}_{\Lambda}$).
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $R_{2}>R_{1}>R_{0}(\Omega)$ and any $k\in\mathbb{N}$,
$\lambda_{k}\left(A^{\mathrm{D}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right).$
@@
where in the last step we used the Cauchy—Schwarz inequality, and the fact that $\lambda_{m}\left(A^{\mathrm{D}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)$ for $1\leq m\leq k$. Hence, with $R_{*}=2R^{*}$, this completes the proof. ∎
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. Then, for every $k\in\mathbb{N}$,
@@
We assume, without loss of generality, that $\int_{\Omega^{\mathrm{ext}}}|\nabla f_{k}|^{2}\,\mathrm{d}x=1$. For any $R>R_{0}(\Omega)$, let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ (respectively its extension by zero outside $\Omega^{\mathrm{ext}}_{R}$), normalised to
$\int_{\Omega^{\mathrm{ext}}_{R}}|\nabla u_{k,R}|^{2}\,\mathrm{d}x=\int_{\Omega^{\mathrm{ext}}}|\nabla{u_{k,R}}|^{2}\,\mathrm{d}x=1.$
@@
Hence, the norm $\|{u_{k,R}}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$ is bounded. Thus, there exists a weakly convergent subsequence in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$, which we again denote by $u_{k,R}$, with weak limit $\widehat{u}_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. The compactness of the trace operator $\boldsymbol{\tau}:E^{1}(\Omega)\to L^{2}(\partial\Omega)$ (see [AuHa14b, Corollary 3.4]) and the convergence of the eigenvalues imply that the convergence of $u_{k,R}$} is strong. Thus, $\widehat{u}_{k}$ is a Steklov eigenfunction with associated eigenvalue $\lambda_{k}\left(A^{\mathrm{D}}\right)$.

### v2

@@ 3.1.2 Weak formulation of the Steklov problem
The following theorem follows from the characterisation [ArtE15, Proposition 5.1] of the operator domain $\operatorname{Dom}\left(A^{\mathrm{D}}\right)$ and [AuHa14b, Section 9].
@@
Additionally, the corresponding eigenfunctions $u_{k}$ can be chosen to form a complete orthonormal basis of the subspace $\left\{u\in E^{1}\left(\Omega^{\mathrm{ext}}\right):\Delta u=0\right\}$ and the restrictions of these eigenfunctions to $\partial\Omega$ form an orthonormal basis of $L^{2}(\partial\Omega)$.
@@
Arendt and ter Elst [ArtE15] proposed two ways to approximate the Dirichlet-to-Neumann operator on exterior domains by truncated domains. For any $R>R_{0}(\Omega)$, recalling that $\Omega^{\mathrm{ext}}_{R}=\Omega^{\mathrm{ext}}\cap B_{R}$, we consider the Dirichlet approximation,
@@
$\begin{cases}\Delta u_{R}=0\qquad&\text{ in }\Omega^{\mathrm{ext}}_{R},\\ \partial_{\nu}u_{R}=\sigma^{\mathrm{N}}u_{R}\qquad&\text{ on }\partial\Omega,\\ \partial_{\nu}u_{R}=0\qquad&\text{ on }\partial B_{R},\end{cases}$ ($\mathrm{SN}_{R}$)
@@
For dimensions $n\geq 3$, it was already proved in [ArtE15] that the truncated domain approaches are equivalent to the approaches from §3.1.1.
Theorem 3.9 ([ArtE15, Theorems 5.5 and 5.6]).
@@
where $\mu=\mu\left(\Omega^{\mathrm{ext}},\Lambda\right)$ is the spectral parameter. Since $\Lambda^{2}>0$, the corresponding bilinear form is $H^{1}\left(\Omega^{\mathrm{ext}}\right)$-coercive. In [Au12, Sections 4 and 8], it is demonstrated that coercivity can be used to show that the spectrum of a Steklov problem on bounded domains is purely discrete. Similarly, by applying the same modifications to [Au12] as in, e.g., [AuHa13, Theorem 5.1], we deduce that the spectrum of ($\mathrm{H}_{\Lambda}$) is purely discrete, with eigenvalues given by
$\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)=\min_{\begin{subarray}{c}M\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)\\ \dim(M)=k\end{subarray}}\max_{\begin{subarray}{c}u\in M\\ u\neq 0\end{subarray}}\frac{\int_{\Omega^{\mathrm{ext}}}|\nabla u|^{2}\,\mathrm{d}x+\Lambda^{2}\int_{\Omega^{\mathrm{ext}}}|u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$ (3.8)
The corresponding eigenfunctions belong to $H^{1}\left(\Omega^{\mathrm{ext}}\right)$, and, in addition, all eigenfunctions are exponentially decaying as $|x|\to\infty$ due to
@@
We now consider the cases of balls of dimension $n\geq 3$ and disks in the planar case $n=2$ separately.
@@
For $n=2$, we have the same form of the eigenfunctions as in (3.9). For $\ell\geq 1$, we proceed as before. For $\ell=0$, we use additionally $K_{0}(x)=-\log\frac{x}{2}+O(1)$ as $x\searrow 0$. This yields
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. We assume that the origin $0\in\Omega$, which can always be achieved by a change of variables, and we identify $\mathbb{R}^{2}$ with $\mathbb{C}$ by interpreting $x=\left(x_{1},x_{2}\right)\in\mathbb{R}^{2}$ as $z=x_{1}+\mathrm{i}x_{2}\in\mathbb{C}$.
@@
We will show later that (3.11) allows us to adapt classical arguments by Weinstock to derive isoperimetric-type inequalities for the first non-trivial Steklov eigenvalue. Prior to that, we will establish, in §5.2, the equivalence between the formulation ($\mathrm{CT}$) and the approaches described in §3.2 and §3.3.
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 2$, is a bounded open set with a Lipschitz boundary $\partial\Omega$. Let $\Phi$ be the fundamental solution for the Laplace equation,
@@
Finally, let us recall that the single layer operator is always invertible in dimensions $n\geq 3$, whereas for $n=2$ we have to assume a suitable scaling of the domain $\Omega$ [St07, Section 6.6]. If $n=2$ and the logarithmic capacity of $\Omega$ is equal to one, the single layer potential has a nontrivial kernel and one cannot use (4.1) to define the eigenvalues of the Dirichlet-to-Neumann operator.
@@
While the boundary integral equation (4.1) remains valid for characterising those eigenfunctions whose harmonic extensions vanish at infinity (up to switching the sign), see §4.3.1, the situation differs in dimension two: as noted in §2.1, exterior Steklov eigenfunctions in two dimensions are only required to remain bounded and do not need to decay. The Dirichlet-to-Neumann operator in exterior domains, including dimension two, together with its Fredholm properties, has been analysed in [KacNaPa25, Section 3]. In contrast to [KacNaPa25], our formulation also incorporates the limit at infinity via layer potentials.
@@
Let $u\in H^{\frac{1}{2}}(\partial\Omega)$, and let $U=\mathcal{H}^{\mathrm{ext}}u\in H^{1}_{\mathrm{loc}}\left(\Omega^{\mathrm{ext}}\right)$ be the harmonic extension of $u$ satisfying (1.2). Green’s identities on $\Omega^{\mathrm{ext}}_{R}$ yield
@@
Additionally, we define the functional $\operatorname{K}_{0}:H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$, as
@@
Accordingly, $\mathcal{D}^{\mathrm{ext}}u=\tau u$ is equivalent to
$\frac{u}{2}-\operatorname{K}_{0}(u)+(\operatorname{K}u)=\tau(\operatorname{V}_{0}u),\qquad u\in H^{\frac{1}{2}}(\partial\Omega).$ ($\mathrm{LP}_{2}$)
@@
where $\left(\cdot,\cdot\right)_{\partial\Omega}$ means the duality pairing $H^{-\frac{1}{2}}(\partial\Omega)\times H^{\frac{1}{2}}(\partial\Omega)\to\mathbb{R}$. Let $w_{\mathrm{eq}}\in H^{-\frac{1}{2}}(\partial\Omega)$ be the unique function satisfying
@@
$w=w_{*}+\alpha w_{\mathrm{eq}},\qquad\text{with}\quad w_{*}\in H_{*}^{-\frac{1}{2}}(\partial\Omega),\quad\alpha=\left(w,1\right)_{\partial\Omega}\in\mathbb{R}.$
@@
$\operatorname{W}(w_{\mathrm{eq}})=-\operatorname{V}(w_{\mathrm{eq}})(0)-\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}\left(w_{\mathrm{eq}},1\right)_{\partial\Omega}=-\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}-\frac{\log(\operatorname{Cap}_{\log}(\Omega))}{2\pi}=0.$
Thus, if $\operatorname{V}_{0}(\alpha w_{\mathrm{eq}})=0$, then $\operatorname{V}(\alpha w_{\mathrm{eq}})=0$ and therefore $\alpha=0$ because $\operatorname{V}$ is invertible. Thus, $\operatorname{V}_{0}$ is injective.
On the other hand, for any $v\in H^{\frac{1}{2}}(\partial\Omega)$ there exists a $w\in H^{\frac{1}{2}}(\partial\Omega)$ with $\operatorname{V}(w)=v$ because $\operatorname{V}$ is invertible. Then, since $\operatorname{W}\left(w\right)\in\mathbb{R}$ and $\operatorname{W}\left(w_{\mathrm{eq}}\right)=0$,
$\begin{split}\operatorname{V}_{0}\left(w+\frac{2\pi\operatorname{W}(w)}{\log(\operatorname{Cap}_{\log}(\Omega))}w_{\mathrm{eq}}\right)&=\operatorname{V}(w)+\frac{2\pi\operatorname{W}(w)}{\log(\operatorname{Cap}_{\log}(\Omega))}\operatorname{V}\left(w_{\mathrm{eq}}\right)+\operatorname{W}(w)\\ &=v+\frac{2\pi\operatorname{W}(w)}{\log(\operatorname{Cap}_{\log}(\Omega))}\frac{-\log(\operatorname{Cap}_{\mathrm{log}}(\Omega))}{2\pi}+\operatorname{W}(w)=v.\end{split}$
Thus, $\operatorname{V}_{0}$ is surjective. ∎
@@ 4.3.1 Decay at infinity
$\frac{1}{2}u+\operatorname{K}u=\tau\operatorname{V}u,\qquad u\in H^{\frac{1}{2}}(\partial\Omega)$ ($\mathrm{LP}_{n}$)
@@
Since $\Phi(x,y)=\frac{1}{(n-2)|\mathbb{S}^{n-1}||x-y|^{n-2}}$, it is easy to see that $U(x)$ satisfies the decay condition (1.2).
@@ 4.3.2 Vanishing flow at infinity
Let $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, be a bounded open set with Lipschitz boundary and $u\in H^{1}_{\text{loc}}(\Omega^{\mathrm{ext}})$ a harmonic function satisfying (4.9). First, note that (4.9) implies that $|\nabla u|\in L^{2}\left(\Omega^{\mathrm{ext}}\right)$. Hence, the average of $u$ as defined in (3.1) and denoted by $\langle u\rangle$, is finite. Define $w:=u-\langle u\rangle$. Then, by definition, $w\in W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$ (see §3.1.1) and therefore, by (3.4), we have $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. By Proposition 6.1, $w=O(|x|^{2-n})$ which shows that a higher-dimensional analogue the of far-field condition (4.3) is satisfied with $u_{\infty}:=\langle u\rangle$.
@@
$\begin{cases}\Delta g=0\qquad&\text{ in }\Omega^{\mathrm{ext}},\\ g=-\frac{1}{\operatorname{Cap}(\Omega)}-\frac{1}{|\mathbb{S}^{n-1}|(2-n)|y|^{n-2}}\qquad&\text{ on }\partial\Omega,\\ g(y)=O\left(|y|^{1-n}\right)\qquad&\text{ as }|y|\to\infty.\end{cases}$
@@
$\begin{split}u_{\infty}=\int_{\partial\Omega}u(y)(-\partial_{\nu_{y}}G_{\mathrm{D}})(y)\,\mathrm{d}S_{y}&=\int_{\partial\Omega}-u(y)\frac{\langle y,\nu_{y}\rangle}{|\mathbb{S}^{n-1}||y|^{n}}\,\mathrm{d}S_{y}-\int_{\partial\Omega}g(y)(\partial_{\nu_{y}}u)(y)\,\mathrm{d}S_{y}\\ &-\int_{\partial B_{R}}\left(g(y)(\partial_{\nu_{y}}u)(y)-u(y)(\partial_{\nu_{y}}g)(y)\right)\,\mathrm{d}S_{y}\end{split}$
@@ 5.1.1 Finite-energy functions and Dirichlet truncation
One could use the resolvent convergence of $A_{R}^{\mathrm{D}}$, shown in [ArtE15], to infer that the spectral problems ($\mathrm{ES}$) and ($\mathrm{SD}_{R}$) yield the same spectrum as $R\to\infty$. Note that the Neumann formulation ($\mathrm{SN}_{R}$) leads to a different spectrum. We instead proceed directly via the variational characterisation. This approach yields a more direct and unified framework which is also applicable to  ($\mathrm{H}_{\Lambda}$).
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. For any $R_{2}>R_{1}>R_{0}(\Omega)$ and any $k\in\mathbb{N}$,
$\lambda_{k}\left(A^{\mathrm{D}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{2}}\right)\leq\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{1}}\right).$
@@
where in the last step we used the Cauchy–Schwarz inequality, and the fact that $\lambda_{m}\left(A^{\mathrm{D}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)$ for $1\leq m\leq k$. Hence, with $R_{*}=2R^{*}$, this completes the proof. ∎
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. Then, for every $k\in\mathbb{N}$,
@@
We assume, without loss of generality, that $\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla{f}_{j},\nabla{f}_{k}\right\rangle\,\mathrm{d}x=\delta_{j,k}$ (where $\delta_{j,k}$ is the Kronecker delta). For any $R>R_{0}(\Omega)$, let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ (respectively its extension by zero outside $\Omega^{\mathrm{ext}}_{R}$), normalised by
$\int_{\Omega^{\mathrm{ext}}_{R}}\left\langle\nabla{u}_{j,R},\nabla{u}_{k,R}\right\rangle\,\mathrm{d}x=\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla{u}_{j,R},\nabla{u}_{k,R}\right\rangle\,\mathrm{d}x=\delta_{j,k}.$
@@
Hence, the norm $\|{u_{k,R}}\|_{E^{1}\left(\Omega^{\mathrm{ext}}\right)}$ is bounded. Thus, there exists a weakly convergent subsequence in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$, which we again denote by $u_{k,R}$, with weak limit $\widehat{u}_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. The compactness of the trace operator $\boldsymbol{\tau}:E^{1}(\Omega)\to L^{2}(\partial\Omega)$ (see [AuHa14b, Corollary 3.4]) and the convergence of the eigenvalues imply that the convergence of $u_{k,R}$} is strong. Hence, for any $v\in E^{1}(\Omega^{\mathrm{ext}})$,
$\begin{split}\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla\widehat{u}_{k},\nabla v\right\rangle\,\mathrm{d}x&=\lim_{R\to\infty}\int_{\Omega^{\mathrm{ext}}}\left\langle\nabla{u}_{k,R},\nabla v\right\rangle\,\mathrm{d}x\\ &=\lim_{R\to\infty}\sigma_{k}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R}\right)\int_{\partial\Omega}{u}_{k,R}v\,\mathrm{d}S=\lambda_{k}\left(A^{\mathrm{D}}\right)\int_{\partial\Omega}\widehat{u}_{k}v\,\mathrm{d}S.\end{split}$
Thus, $\widehat{u}_{k}$ is a Steklov eigenfunction with associated eigenvalue $\lambda_{k}\left(A^{\mathrm{D}}\right)$.